OpenStudy (anonymous):

Is this a quadratic equation? And how do i solve if it is? y = x2 − 4x

OpenStudy (anonymous):

A quadratic equation is one in the form of: y=Ax^2 +Bx + C. Given the equation you have provided, it is a quadratic. That is, for y=x^2 - 4x, A=1, B=-4 and C=0. Basically what I'm trying to say here is : it is a quadratic. Now by solve, I'm going to assume you're asking how to find it's x-intercepts. We do this by letting y=0. That is: x^2 -4x=0 -> x(x-4)=0 Hence the values of x which satisfy this equation are x=0 and x=4.

OpenStudy (anonymous):

x or (x-4) =0 so values are 0 or 4

OpenStudy (anonymous):

what is the vertex of it then?

OpenStudy (anonymous):

@jimswig88 @extremity

OpenStudy (anonymous):

 is that how it should look to solve the vertex then?